# DeepQuestion: Systematic Generation of Real-World Challenges for Evaluating LLMs Performance

### v1

@@ Abstract
LLMs often excel on standard benchmarks but falter on real-world tasks. We introduce DeepQuestion, a scalable automated framework that augments existing datasets based on Bloom’s taxonomy and creates novel questions that trace original solution paths to probe evaluative and creative skills. Extensive experiments across ten open-source and proprietary models—covering both general-purpose and reasoning LLMs—reveal substantial performance drops (even up to 70% accuracy loss) on higher-order tasks, underscoring persistent gaps in deep reasoning. Our work highlights the need for cognitively diverse benchmarks to advance LLM progress. DeepQuestion and related datasets will be released upon acceptance of the paper.
@@
Ali Khoramfar, Ali Ramezani, Mohammad Mahdi Mohajeri Mohammad Javad Dousti, Majid Nili Ahmadabadi, Heshaam Faili Department of Electrical and Computer Engineering University of Tehran {khoramfar, ali.ramezani.96, mehdimohajeri, mjdousti, mnili, hfaili}@ut.ac.ir
1 Introduction
Large Language Models (LLMs) have shown strong performance on a variety of standardized benchmarks, such as GSM8K, MATH, MMLU, MMLU-Pro, and GPQA Achiam et al. (2023); Team et al. (2023); Grattafiori et al. (2024); Team et al. (2025); Cobbe et al. (2021); Hendrycks et al. (2021b, a); Wang et al. (2024); Rein et al. (2024). However, these benchmarks are highly structured and may not capture the complexity and variability found in real-world problems Arias-Duart et al. (2025); Myrzakhan et al. (2024).
Recent work GSM-Symbolic Mirzadeh et al. (2025) shows that adding extra numerical values and irrelevant information to benchmark questions reduces LLM performance, highlighting their difficulty with extraneous content. Since real-world tasks often include such irrelevant details, this limitation raises concerns about the practical robustness of LLMs.
Proof or Bluff Petrov et al. (2025) finds that while LLMs perform well on older math contests, they struggle with the 2025 USA Mathematical Olympiad, with top models like Gemini-2.5-Pro scoring below 25. This indicates a reliance on training data patterns rather than genuine reasoning abilities.
Alaa et al. (2025) report that while LLMs excel on medical exams like the USMLE, these tests lack construct validity and do not reflect real clinical complexity. Thus, strong benchmark results do not guarantee clinical effectiveness. Scenario-based questions that assess reasoning, rather than recall, are needed for a more accurate evaluation of higher-order cognition Chauhan et al. (2025).
Learning science offers frameworks for understanding learning levels, Bloom’s Taxonomy Forehand (2010) is a widely recognized framework outlining six cognitive skill levels, from basic recall to higher-order reasoning Smith (2008); Leung (2000). LLMs often excel at recall tasks but struggle with novel problems Petrov et al. (2025), indicating limited generalization beyond memorized data.
Building on Bloom’s Taxonomy, we introduce the DeepQuestion framework to create new benchmarks from existing datasets. It generates (1) real-world scenarios to test application and analysis, and (2) instructions for crafting similar questions to assess creation, evaluation, and instruction-following, enabling deeper evaluation of LLMs’ knowledge. Our main contributions are:
@@ 1 Introduction
The remainder of this paper is organized as follows: Section 2 introduces the DeepQuestion framework, Section 3 details our experimental setup, and the final sections present our results and conclusions.
2 Method
@@
This study is guided by a key research question: Can existing benchmarks be deepened to reflect real-world tasks more accurately?
To address this question, we draw on educational learning science, especially Bloom’s taxonomy, which organizes cognitive skills into six levels. Prior works show that most benchmark questions focus on the first two levels—remembering and understanding Alaa et al. (2025); Chauhan et al. (2025). We hypothesize that deepening these questions along Bloom’s hierarchy could better engage higher-order reasoning and reflect real-world complexity. Therefore, we introduce the DeepQuestion framework, which generates more insightful evaluation questions through two methods: (1) question-to-scenario (Q2S) and (2) question-to-instruction (Q2I) generation. The details of these methods, as well as the prompt generation pipeline that automates their application to existing benchmarks, are described in the rest of this section.
2.1 Question-to-Scenario Generation (Q2S)
@@
To deepen questions, Q2S targets the third level of Bloom’s taxonomy, namely apply, which emphasizes on using knowledge in practical contexts. Each original question is transformed into a scenario-based version embedding the core problem within a realistic narrative, often including extraneous details and some distractions to simulate real-world complexity. For example, a basic physics question asking for the position-time equation of an object, is reframed within a scenario involving an environmental drone, where irrelevant elements like weather acted as distractors. The core question data and required reasoning remain unchanged. Q2S employs an automated prompt-based pipeline powered by an LLM (see Subsection 2.3 for more details on the pipeline). The output is a benchmark of scenario-driven questions that better reflect authentic problem-solving contexts.
2.2 Question-to-Instruction Generation (Q2I)
@@
Encouraged by the initial result analysis at the Bloom’s apply level with Q2S, we extend our approach to the higher levels of evaluate and creation. At these levels, learners are expected to assess information critically and produce novel outputs. We hypothesize that individuals or models with deep conceptual understanding are not only able to solve problems but also to design meaningful questions. This informs our second experimental setup, which reverses the direction of the task: instead of solving a question, the model is instructed to design one.
@@
For each original dataset question, Q2I constructs an instruction designed to prompt a language model to generate a new question that preserves the same topic and solution path as the original. This instruction is formulated in a way that ensures a different LLM, when given the instruction, would also produce a question aligned with the original in both topic and reasoning process. For example, if the original question in Subsection 2.1 (position-time equation of an object) yields the solution $x=3t-4$, the instruction would ask the model to design a physics problem involving uniform linear motion whose solution would be exactly that equation. Designing questions based on instruction implicitly tests three levels of the model’s reasoning: conceptual understanding of the domain and relevant equations, the ability to define appropriate variables and their interrelations, and the selection of values that yield the target solution.
@@
2.3 Prompt Generation Pipeline
Since transforming each benchmark into its Q2S and Q2I variants requires carefully crafted prompts tailored to the specific domain and style of that benchmark, a key component of the DeepQuestion framework is a prompt generation pipeline which automates the creation of these task-specific prompts and is shown in Figure 1. This pipeline eliminates the dependency on human expertise for prompt design and enables scalable adaptation to different benchmarks.
@@ 2.3 Prompt Generation Pipeline
If the score surpasses a threshold (e.g., 8), the prompt is accepted for use in the corresponding generation task. Otherwise, the prompt generator revises the prompt using the evaluator’s feedback, and the process repeats iteratively until a satisfactory prompt is produced.
This automatic prompt optimization is described in Algorithm 1 in the Appendix Section. By leveraging this pipeline, the DeepQuestion framework achieves robust prompt design tailored to the specific domain and style of each benchmark, enabling effective question transformation without manual intervention.
3 Experiment Setup
(a) Original vs. Q2I in GSM8K
(b) Original vs. Q2I in Physics
Figure 2: Evaluation of different LLMs in original and instruction-based question
In the Q2I setting, the model generates questions from instructions. Evaluation consisted of two steps: (1) Answerability Check: verifying whether a strong model (O4-mini) and the model itself can correctly solve the generated questions, indicating questions validity. (2) LLM-as-Judge: since answerability alone does not capture question quality, a powerful language model (O4-mini) was employed for direct qualitative evaluation based on expert-defined criteria.
4 Results
4.1 General-Purpose LLMs
First, we evaluated general-purpose LLMs on question generation using Q2S. Moderate performance loss in Q2S (see Appendix C) motivated us to design the Q2I experiment, where models showed a substantial accuracy drop for tasks demanding deeper understanding. No model surpassed 38% accuracy on instruction-based question generation for GSM8K or physics (see Figure 2), despite over 95% on original questions. These results support the use of Bloom’s taxonomy as an evaluation framework and reveal the limitations of general-purpose LLMs on higher-order cognitive tasks.
4.2 Reasoning LLMs
Reasoning models maintain relatively stable performance in the Q2S setting, indicating proficiency at the Bloom’s apply level (Appendix C). However, in the Q2I setting, accuracy drops by 8–30% (Figure 2), exposing limitations at higher cognitive levels. For instance, Deepseek-R1’s performance decreases by 30% on GSM8K and 10% on physics, while Qwen3-32b drops by 11% on both. Although reasoning models outperform general-purpose models, substantial gaps remain in achieving deep, comprehensive understanding.
Figure 3: Win rate of original against Q2I physics questions for O4-mini.
4.3 Question Quality Analysis
While reasoning models can generate questions that follow instructions and maintain topical consistency, it is crucial to assess their quality beyond surface-level compliance. To this end, we use a preference-based evaluation with O4-mini as a judge, directly comparing generated and original questions across five criteria: reasoning demand, numerical quality, physical realism, clarity and brevity, and solution spoiling (see Appendix B for explanation of each criterion). This approach addresses concerns about the model’s ability to assess quality by focusing on direct comparison.
As shown in Figure 3 and Appendix D, generated physics questions still lag behind the originals, reflecting ongoing challenges for reasoning models. In contrast, generated GSM8K questions are more comparable to the originals, likely due to the dataset’s simplicity. These results confirm the effectiveness of our framework in increasing question depth without introducing artificial difficulty. However, a significant quality gap persists, highlighting the ongoing challenge of generating high-quality questions, even for advanced reasoning models.
5 Conclusion
Despite strong performance on standard benchmarks, LLMs struggle with tasks requiring deeper and real-world reasoning. We introduced DeepQuestion, a scalable automated framework which enhances existing datasets based on Bloom’s taxonomy. Our method exposes reasoning gaps often missed by conventional evaluations. Applied to GSM8K and physics datasets, our framework reveals sharp accuracy drops, up to 70%, on complex reasoning tasks across a wide range of models. These findings highlight serious limitations in current LLMs and underscore the need for cognitively rich benchmarks.
Limitations
Although our proposed framework is capable of handling open-ended questions, in this study we restrict our focus to questions with definite correct answers. To enable expert human evaluation, we concentrate on two benchmarks related to mathematics and physics. Nonetheless, our framework is not specifically tailored to these benchmarks and can be extended to other domains as well.
Appendix A Algorithm
Algorithm 1 describes the iterative dual-LLM pipeline for generating and evaluating prompts that transform an existing QA dataset into deep, scenario- and instruction-based questions.
@@ 9      $(score,feedback)\leftarrow\texttt{LLM}(evaluation\_criteria,prompt)$;
13       $q^{\prime}\leftarrow\texttt{LLM}(prompt,q,a)$; DeepQuestions.append$(q^{\prime},a)$;
@@
Appendix B Prompts
Below are the core prompt templates used in the DEEPQUESTION framework, including the Q2S scenario-generation and Q2I instruction-generation prompts, as well as the system-level prompts for prompt evaluation and question generation.
System Prompt for Questions Quality Evaluation in Q2I You are an expert evaluator of high school physics problems. You will compare two problems, A and B, which lead to the same numeric answer (provided only for verification). Your job is to judge which problem is better, based on meaningful teaching and assessment value — not on how easy or flashy it appears. Before making your judgment, consider: 1. What physical reasoning each problem requires. 2. How the numbers and context affect realism and difficulty. 3. How clearly and economically the problem is written. 4. Whether the problem gives away the solution or answer inappropriately. — ### PROBLEMS A: {original_q} B: {generated_q} target_answer = {target_answer} — ### CRITERIA & GUIDANCE 1. Reasoning Demand - Definition: How much actual thinking and decision-making the problem requires. - Better if: The solver must interpret the situation, choose a method, or justify an approach — not just apply a memorized formula. - Worse if: The solution is a single plug-in step with no real choices or insight required. 2. Numerical Quality - Definition: Thoughtfulness and realism of the numbers used. - Better if: Values are plausible and require genuine calculation; they are not trivial. - Worse if: The numbers oversimplify the math or seem artificially chosen just to make the answer neat (e.g., all variables are $1$ or 00). 3. Physical Realism - Definition: Whether the scenario is physically plausible and internally consistent. - Better if: Units, magnitudes, and assumptions make sense; any simplifications (e.g., ignoring friction) are clearly stated. - Worse if: The context contradicts basic physics or feels contrived or incoherent. 4. Clarity & Brevity - Definition: How concise and understandable the problem is. - Better if: The question is short, clear, and includes only relevant information. - Worse if: It contains unnecessary explanation, formula reminders, solution hints, or redundant language. 5. Solution Spoiling - Definition: Whether the problem gives away the answer or guides through steps. - Better if: The question asks for a result without hinting at how to obtain it. - Worse if: The final answer is stated, or if key solution steps are included (e.g., "first calculate time, then…"). — ### DECISION RULES - For each criterion, assign the better problem: A, B, or E (Equal). - Use E only if the difference is truly negligible — avoid it unless absolutely necessary. - Be decisive and clear. — ### OUTPUT Produce exactly this JSON (no extra text): REMEMBER: use E only when A and B are nearly indistinguishable. {{ "Reasoning_Demand": "<A|B|E>", "Numerical_Quality": "<A|B|E>", "Physical_Realism": "<A|B|E>", "Clarity & Brevity": "<A|B|E>", "Solution_Spoiling": "<A|B|E>" }}

### v2

# arXiv:2505.24532v2[cs.CL] 27 Feb 2026

@@ Abstract
While Large Language Models (LLMs) achieve near-human performance on standard benchmarks, their capabilities often fail to generalize to complex, real-world problems. To bridge this gap, we introduce DeepQuestion, a scalable, automated framework that systematically elevates the cognitive complexity of existing datasets. Grounded in Bloom’s taxonomy, DeepQuestion generates (1) scenario-based problems to test the application of knowledge in noisy, realistic contexts, and (2) instruction-based prompts that require models to create new questions from a given solution path, assessing synthesis and evaluation skills. Our extensive evaluation across ten leading open-source and proprietary models reveals a stark performance decline—with accuracy dropping by up to 70%—as tasks ascend the cognitive hierarchy. These findings underscore that current benchmarks overestimate true reasoning abilities and highlight the critical need for cognitively diverse evaluations to guide future LLM development.
Keywords: Evaluation Methodologies, Cognitive Methods, Learning Science inspired Evaluation
[…]
@@
Ali Khoramfar, Ali Ramezani, Mohammad Mahdi Mohajeri
Mohammad Javad Dousti, Majid Nili Ahmadabadi, Heshaam Faili
Department of Electrical and Computer Engineering
University of Tehran
{khoramfar, ali.ramezani.96, mehdimohajeri, mjdousti, mnili, hfaili}@ut.ac.ir
Abstract content
1.  Introduction
Recent advances in large language models (LLMs) have driven remarkable improvements across a wide spectrum of benchmarks, from arithmetic reasoning in GSM8K to expert-level performance on MMLU and GPQA Achiam et al. (2023); Team et al. (2023); Grattafiori et al. (2024); Team et al. (2025); Cobbe et al. (2021); Hendrycks et al. (2021b, a); Wang et al. (2024); Rein et al. (2024).. Yet, as scores approach saturation, evidence increasingly suggests that such benchmarks do not reflect the demands of real-world reasoning. When confronted with tasks that include irrelevant details, ambiguous contexts, or creative synthesis, even state-of-the-art LLMs exhibit significant failures—revealing an overreliance on pattern recognition rather than genuine understanding Arias-Duart et al. (2025); Myrzakhan et al. (2024).
Emerging studies have begun to expose this gap. Models that perform flawlessly on structured academic datasets often falter on problems drawn from authentic or modern settings, such as new mathematics contests, clinical scenarios, or symbolic reasoning variations Shojaee et al. (2025); Mirzadeh et al. (2025); Petrov et al. (2025); Alaa et al. (2025). These shortcomings highlight that current benchmarks primarily target lower-order cognitive abilities—such as recall and comprehension—while neglecting the deeper reasoning, analysis, and creativity required for practical decision-making. Without cognitively diverse evaluation frameworks, progress risks being misinterpreted as intelligence rather than memorization.
To address this limitation, we introduce DeepQuestion, a systematic framework for generating cognitively enriched evaluation datasets grounded in Bloom’s taxonomy of learning Forehand (2010); Leung (2000) that is a widely recognized framework outlining six cognitive skill levels, from basic recall to higher-order reasoning (see Figure˜1). DeepQuestion transforms existing question–answer pairs into (1) scenario-based problems that test the application of knowledge in realistic, constraint-rich situations, and (2) instruction-based prompts that assess creation and evaluation skills by requiring models to design new questions aligned with given solution paths. This taxonomy-guided approach adds interpretable layers of complexity that correspond directly to human cognitive processes.
Figure 1: Bloom’s taxonomy hierarchic
Through extensive experiments across ten leading LLMs—spanning general-purpose and reasoning-oriented architectures—we observe sharp performance declines, up to 70%, as the cognitive level of tasks increases. These results provide evidence that LLMs’ strong benchmark performance does not generalize to deeper reasoning. Beyond evaluation, the DeepQuestion framework offers a replicable foundation for constructing cognitively meaningful benchmarks across diverse disciplines, from mathematics and physics to law and medicine. Our main contributions are:
@@ 1.  Introduction
The remainder of this paper is organized as follows: Section˜2 reviews the related works, Section˜3 introduces the DeepQuestion framework, Section˜4 details our experimental setup, and the final sections present our results and conclusions.
2.  Related Works
This paper draws on ideas from two main areas: the levels of learning, derived from the learning sciences, and the limitations of LLMs in solving complex and real-world problems. In this section, we review the previous work related to these two domains.
2.1.  Learning Levels
In the learning sciences, various definitions of learning have been proposed. Some of these frameworks conceptualize learning as occurring at different levels. Among the most influential are the SOLO taxonomy and Bloom’s taxonomy, both of which define learning across hierarchical stages İlhan et al. (2017).
In the SOLO taxonomy Biggs and Collis (2014), the lowest level of learning is Pre-structural, where no meaningful learning has occurred. The highest level, known as Extended Abstract, indicates that the learner is able to extend their understanding and apply their knowledge to new domains.
Bloom’s taxonomy Bloom (1971); Anderson and Krathwohl (2001); Forehand (2010) offers a more widely adopted framework. It also defines learning in multiple levels and exists in several versions. In its most recent revision, the lowest level is remember, in which the learner simply recalls information without deeper understanding. A higher level, apply (the third level), represents the learner’s ability to use acquired knowledge to solve real-world problems. The highest level, create, reflects the learner’s capacity to generate new ideas or concepts based on previously acquired knowledge.
Employing these taxonomies to assess the performance of LLMs can provide a more human-aligned perspective on evaluating their knowledge and reasoning capabilities.
$2.$  LLMs Evaluation in Complex and Real-world Problems
With the rapid advancement of LLMs, their ability to solve complex problems has become one of the key questions in the field. Moreover, as the potential use of LLMs in real-world contexts grows, evaluating their performance on real-world problems has become increasingly important.
MMLU Hendrycks et al. (2021a) serves as a standard benchmark for evaluating LLMs across diverse academic and professional domains. However, as model performance on MMLU has approached saturation, MMLU-Pro Wang et al. (2024) was introduced to assess models on more complex, reasoning-intensive tasks. Establishing such advanced benchmarks is essential to accurately measure progress and ensure meaningful evaluation of LLMs’ capabilities in challenging real-world scenarios. This aligns with findings from The Illusion of Thinking Shojaee et al. (2025), which highlight the limitations of current LLMs when confronted with more complex reasoning tasks.
Recent studies Chauhan et al. (2025); Alaa et al. (2025) have examined the performance of LLMs in real-world scenarios and compared it with their performance on standard benchmarks. These studies indicate a notable reduction in LLM performance when applied to real-world contexts. Given the growing interest in deploying these models in practical applications, assessing their performance in real-world settings is crucial. Consequently, rigorous evaluation of LLMs under such conditions has become increasingly important.
Figure 3: Examples of question transformations produced by the Q2S and Q2I pipelines
3.  Method
@@
This study is guided by a key research question: Can existing benchmarks be systematically deepened to reflect real-world task complexity, rather than introducing artificial difficulty?
To address this, we leverage insights from learning science—particularly Bloom’s taxonomy—which offers a principled framework for systematically controlling and measuring cognitive complexity. Unlike ad-hoc increases in difficulty, Bloom’s hierarchy allows us to add real, interpretable complexity to benchmark items, specifying exactly how and by how much each transformation increases the cognitive demands.
Prior analyses reveal that most benchmarks remain at lower Bloom levels—remember and understand Alaa et al. (2025); Chauhan et al. (2025)—and rarely stress the higher-order reasoning required in practical tasks. To bridge this gap, we focus on the apply and create levels, which represent critical forms of real-world reasoning often underrepresented in evaluation. Specifically, we propose two aligned methods: (1) question-to-scenario (Q2S), which increases complexity by embedding items in realistic, constraint-rich contexts to elicit apply-level reasoning; and (2) question-to-instruction (Q2I), which reframes items as multi-step, constructive procedures with explicit acceptance criteria, targeting the create level of synthesis and innovation. These transformations go beyond surface-level difficulty and require test-takers to transfer, adapt, and synthesize knowledge in meaningful, task-driven ways.
By systematically deepening benchmarks along Bloom’s hierarchy using Q2S and Q2I, our DeepQuestion framework enables precise diagnosis of model weaknesses—revealing, for example, whether a model fails at context adaptation, constraint handling, or constructive synthesis. This fine-grained understanding directly informs concrete intervention strategies for model improvement, such as targeted data augmentation, objective design, or architectural changes. Implementation details and our automated prompt-generation pipeline are described in the following section.
3.1.  Question-to-Scenario Generation (Q2S)
@@
To deepen questions, Q2S targets the third level of Bloom’s taxonomy, namely apply, which emphasizes on using knowledge in practical contexts. Each original question is transformed into a scenario-based version embedding the core problem within a realistic narrative, often including extraneous details and some distractions to simulate real-world complexity. For example, as shown in Figure˜3, a basic physics question asking for the position-time equation of an object, is reframed within a scenario involving an environmental drone, where irrelevant elements like weather acted as distractors. The core question data and required reasoning remain unchanged. Q2S employs an automated prompt-based pipeline powered by an LLM (see Subsection˜3.3 for more details on the pipeline). The output is a benchmark of scenario-driven questions that better reflect authentic problem-solving contexts.
These scenario-based questions resemble real-world problems. To solve them, LLMs must extract relevant information from the scenario and apply their knowledge accordingly. This process reflects a higher cognitive level than the stages of remember and understand.
3.2.  Question-to-Instruction Generation (Q2I)
@@
Encouraged by the initial result analysis at the Bloom’s apply level with Q2S, we extend our approach to the higher levels of evaluate and create. At these levels, learners are expected to assess information critically and produce novel outputs. We hypothesize that individuals or models with deep conceptual understanding are not only able to solve problems but also to design meaningful questions. Designing a question requires the evaluation of existing knowledge in order to shape a coherent concept. It also involves the ability to establish a logical and structured process that leads to the formulation of a meaningful question. This process demands both critical reflection and creative synthesis of ideas. Consequently, the capacity to design questions can be used to assess competencies at both the evaluate and create levels of cognition. This informs our second experimental setup, which reverses the direction of the task: instead of solving a question, the model is instructed to design one.
@@
For each original dataset question, Q2I constructs an instruction designed to prompt an LLM to generate a new question that preserves the same topic and solution path as the original. This instruction is formulated in a way that ensures a different LLM, when given the instruction, would also produce a question aligned with the original in both topic and reasoning process. For example, if the original question in Figure˜3 (position-time equation of an object) yields the solution $x=5t+t^{2}$, the instruction would ask the model to design a physics problem whose solution would be exactly that equation. Designing questions based on instruction implicitly tests three levels of the model’s reasoning: conceptual understanding of the domain and relevant equations, the ability to define appropriate variables and their interrelations, and the selection of values that yield the target solution.
@@
$3.$  Prompt Generation Pipeline
Since transforming each benchmark into its Q2S and Q2I variants requires carefully crafted prompts tailored to the specific domain and style of that benchmark, a key component of the DeepQuestion framework is a prompt generation pipeline which automates the creation of these task-specific prompts and is shown in Figure˜2. This pipeline eliminates the dependency on human expertise for prompt design and enables scalable adaptation to different benchmarks.
@@
If the score surpasses a threshold (e.g., 8), the prompt is accepted for use in the corresponding generation task. Otherwise, the prompt generator revises the prompt using the evaluator’s feedback, and the process repeats iteratively until a satisfactory prompt is produced. Once a prompt is accepted, it is then reviewed by a human domain expert, who either approves or rejects the generated prompts. In our experiments on the GSM8K and physics question datasets, the produced prompts were approved by the experts. See the generated prompt for Q2S for physics questions in the following.
This automatic prompt optimization is described in Algorithm˜1. By leveraging this pipeline, the DeepQuestion framework achieves robust prompt design tailored to the specific domain and style of each benchmark, enabling effective question transformation without manual intervention.
@@ 9 $(score,feedback)\leftarrow\texttt{LLM}(evaluation\_criteria,prompt)$;
[…]